# ElectricityEmissions.jl: A Framework for the Comparison of Carbon Intensity Signals

### v2

@@ 4.1. Carbon Intensity Calculation
The first part of the code example demonstrates how to compute the carbon intensity metrics. The network data and generator emissions information is loaded using the parse_file function from the PowerModels.jl package. It is passed to the functions calculate_ACE, calculate_LMCE, calculate_LACE and calculate_ALMCE to calculate the respective carbon intensity metrics. Each of these functions solve a piece-wise linear DCOPF problem to dispatch the system. Then load demand, optimal generator set-points, and generatorcarbon intensity (and for some metrics, the network structure) are used to calculate the nodal carbon intensity for each node in the network, as described in Section 3. We calculate total system emissions with the function calculate_system_emissions, which only needs generator set-points as an input.
@@ 4.3. Example Workflow
2. Solve a Load-Shifting Problem Use the carbon intensity values generated in the previous step (or outside carbon metric values/forecasts) as input to solve a load shifting optimization problem. Do this for each intensity metric and record the resulting post-shift loads. To estimate the impact of this shifting on the accounted carbon emissions, multiplying the initial carbon intensities by the new load values. Note that these are just estimates, for two main reasons. First, the carbon intensities might change as a result of the load shifting, leading to differences between the estimated and realized accounted carbon emissions. Second, the carbon intensities are themselves estimates of the carbon footprint of electricity. Although a load shifting action leads to a significant reduction in the accounted carbon emissions, it is not guaranteed that the total system emissions would actually decrease by the same amount.
3. Post-Shift Emissions Calculation Shifting load will result in new generator set-points and a new power flow which may change the carbon intensity values at each bus. To determine the final post-shift emissions, we rerun the relevant calculate_...() functions with the shifted load profile to recalculate the carbon intensity metrics. As explained above, the carbon intensities may change as a result of the load shifting. Using the post-shift carbon intensities, it is possible to compute for the accounted emissions that would be realized if the load-shifting was implemented. We refer to those as the realized accounted emissions.
@@ 5.1. Dataset
We make use of the RTS-GMLC (Barrows et al., 2020) test case, which contains 73 buses, 120 lines, and 158 generators. One year’s worth of hourly load and generation data is available for this system, totaling 8,784 cases. We make two modifications/additions, similar to the changes made to the same system in (Lindberg et al., 2021a). First, we add data-centers to the system by setting buses 103, 107, 204, and 322 to be data-centers, each with a nominal load of 250MW. Second, we use values from the U.S. Department of Energy(Office of Energy Policy and Systems Analysis, U.S. Department of Energy, 2016) to assign the following emissions intensities (in metric tons of CO2 per MWh) to each generator in the system:
@@
The data and code are available as examples in the package documentation  [^2] https://github.com/WISPO-POP/ElectricityEmissions.jl.
@@ 5.3. Results
Carbon Emissions Intensity
(Tons CO2/MWh)
Whole
System
DC 103 DC 107 DC 204 DC 322
Metric Mean SD Mean SD Mean SD Mean SD Mean SD
@@
Table 1. Whole-Year Emissions Intensity Data
Carbon Emissions
(Million Tons CO2/MWh)
True
System
System
Acct.
DC
Acct.
DC 103
Acct.
DC 107
Acct.
DC 204
Acct.
DC 322
Acct.
Metric Sum Sum Sum Sum Sum Sum Sum
LMCE 15.828 33.012 6.692 1.686 1.562 1.917 1.527
ALMCE 15.828 3.162 0.803 0.679 1.035 0.644
ACE 15.828 3.008 0.752 0.752 0.752 0.752
LACE 15.828 2.707 0.577 0.850 1.153 0.126
Table 2. Whole-Year Emissions Accounting Data
@@
We first consider how the choice of carbon intensity metric impacts the total accounted carbon emissions across the year. Table 2 contains the accounted emissions for the entire system and each data-center location, as well as the true system emissions as measured by the generator set-points. In looking at the emissions assigned to the entire system, we see that ALMCE, ACE, and LACE leads to accounted emissions that equal the true system emissions value of 15.8 million tons of CO2. LMCE, however, lead to much higher accounted emissions at 33.0 million tons. This trend of much higher emissions with LMCE also holds true for the different data center locations.
@@
We next examining how the various carbon metrics assign emissions intensities to different system loads. Table 1 shows the emissions intensity statistics across the whole system, and for each data-center load individually. We show both the mean value across all hours of the year, as well as the standard deviation of the emissions intensity. First, we analyze inter-metric intensity differences for the system as a whole. On average, LMCE has the highest intensity values across the whole system, with a mean value of 0.740 Tons CO2/MWh. This is dramatically higher than the next-highest, ACE, with a mean intensity of 0.342 Tons CO2/MWh. ACE shows the least whole-system variability out of all the metrics, likely due to the fact that it does not vary spatially within a given time-step.
Second, we analyze emissions intensities across the four data-centers, located at buses 103, 107, 204, and 322. As in the whole-system statistics, LMCE assigns the highest average intensity values of any metric to all data-centers. However, the relative magnitude of ALMCE, ACE, and LACE varies greatly between location. Out of these three metrics, LACE has the highest mean value for the data-centers at buses 107 and 204, but the lowest for those at buses 103 and 322. Despite this variability, all of the metrics (save ACE, which is location-agnostic) are in agreement on the ranking of average emissions intensity between data-center locations. DC 204 has the highest emissions intensity, followed by DC 103, DC 107, and finally DC 322.
@@ 6.1. Dataset
We make use of the same modified RTS-GMLC test case as described in the accounting-only case study, which provides a year’s worth of hourly load and generation data. Whereas the accounting-only case study added fixed 250MW data-center loads, the data-centers are new allowed to vary their load in order to minimize emissions.
@@ 6.2. Methodology
We perform load shifting based on LMCE, ALMCE, ACE, and LACE carbon metrics, within each 24-hour period of the dataset. Following the workflow described in Section 4.3, we first calculate initial ”pre-shift” emissions intensity for each metric.
Next, the data-centers, which are assumed to operate in a network controlled by a single entity, are then allowed to shift their load both in space and time to align with periods/locations of low carbon intensity. The mathematical formulations of the data center load shifting problem can be found in Appendix C. As an outcome of the optimization problem, we obtain estimated accounted emissions for the data centers.
Finally, the load-shifting step is followed by a re-dispatch of the system and recalculation of carbon intensity metrics in order to determine realized accounted emissions and whole-system emissions outcomes.
@@ 6.3. Results
Carbon Emissions (Million Tons CO2)
Pre-Shift Accounted Emissions
Post-Shift Accounted Emissions
Post-Shift Accounted Emissions
@@
Metric
Whole
System
DC
Loads
Non-DC
Loads
DC
Loads
Whole
System
DC
Loads
Non-DC
Loads
Whole
System
@@
Table 3. Shifting Case Study: Emissions Results
@@
Starting on the left, we show the pre-shift accounted emissions which are the carbon accounting values prior to any shifting (i.e. they are the same values as in the accounting case study, summarized here for convenience of the reader). Post-shift accounted emissions are divided into ”estimated” and ”realized”.
Next we show the estimated post-shift accounted emissions, which refer to the data-centers’ belief about what their accounted emissions will be after shifting, based on their new (post-shift) load values and the old (pre-shift) carbon metric values that they were optimizing for. We only show estimated post-shift accounted emissions for the data centers themselves, as these are the only loads that perform load shifting and obtain such estimates.
The realized post-shift emissions are the actual accounted emissions values, determined via post-shift loading and post-shift (i.e. re-calculated) carbon intensity values.
@@
We begin by investigating whether load-shifting leads to a reduction in the accounted emissions assigned to data-center loads. To analyze this, we first analyze the realized accounted emissions for DC loads. We observe that shifting according to any of the metrics leads to a decrease the accounted emissions for DC loads, but to varying degrees. The greatest reduction in accounted emissions was achieved by ALMCE, folloed by LACE and ACE. The smallest reduction in accounted emissions happened when the data centers shifted according to LMCE.
@@
In conclusion, these results indicate that even though data centers are able to claim emission reductions after shifting with ALMCE, ACE and LACE, these emission reduction come from reallocating emissions to other loads rather than from actually decreasing emissions. In contrast, when shifting with LMCE, the data centers actually reduced total system emissions.
@@
While over the course of the whole year, LMCE/LACE-based shifting seemed to decrease emissions and ALMCE/ACE-based shifting seemed to increase them, there is a large amount of variability in the performance of each metric on any particular day. Figure 4 shows the distribution of the daily changes in system emissions resulting from shifting with the different metrics, as well as the mean, median, 10th percentile and 90th percentile of the reductions. We can clearly see that shifting based on LMCE is much more likely to lead to large carbon emission reductions (i.e. large negative values) and less likely to increase the total carbon emissions from the overall system (i.e. has fewer positive values). We can also clearly see how some of the other metrics have a tendency to increase carbon emissions rather than reducing them.
@@
Finally, we briefly examine the outcome of shifting based on one metric and accounting using another (e.g. determining data-center loads with LMCE, then assigning emissions using ACE). Figure 5 shows the absolute change in assigned data-center emissions, defined as the difference between pre- and post-shift accounted emissions (measured by the accounting metric).
As is to be expected, when the same metric is used for shfiting and accounting, we see the greatest reduction in accounted emissions. There is great variation in the ’cross-metric’ behavior of different ”shifting metric”-”accounting metric” pairs. For many of them, we see that shifting with one metric (e.g. LMCE) increases the accounted carbon emissions when a different metric is used for accounting. This indicates that companies or organizations who want to not only reduce overall carbon emissions, but also the carbon emissions that are allocated to them, have strong incentives to shift according to the same metric that is used for accounting.
@@ 7. Conclusions
In our case study, we find that load shifting based on the ACE, ALMCE and LACE intensity metrics has very little impact on overall carbon emission and instead tends to shift emissions from the data centers (which actively seek to minimize their emissions according to these metrics) to non-data center loads. In contrast, shifting according to LMCE reduces emissions for both data centers, non-data centers and the overall system. However, it is challenging to do accounting with LMCE, as the allocated carbon emissions are more than two times higher than the actual carbon emissions of the system.
@@
In future work, we plan to expand the package to include other metrics and other functionality, allowing input from the community. We also wish to design new metrics with nice properties for both accounting and load shifting, and perform more thorough analysis of the differences between metrics on more realistic power system test cases and more realistic load-shifting setups.
Appendix A DCOPF Formulation
We use the following DCOPF formulation with piecewise linear costs to dispatch generation in ElectricityEmissions.jl, and for the linear generation sensitivity calculations necessary for the LMCE and ALMCE carbon metrics:
@@ Appendix A DCOPF Formulation
Each generator has a $N_{ct}$-term piecewise-linear cost function. Equation 6b constrains the cost of each generator $j$ to be above all of the $k$ linear functions ($f_{j,k}$) that define its segments. Equation 6c is the standard nodal power conservation constraint, referencing generation ($G_{i}$), loads ($D_{i})$, and connected lines $(i,j)\in L$ at bus $i$. Equation 6d symmetrically limits the flow on each line $(i,j)$ by $P_{i,j}^{lim}$, and Equation 6e sets limits for each generator $j$ to be between $P_{G,j}^{min}$ and $P_{G,j}^{max}$. Finally, Equation 6f sets the voltage angle of the reference bus to zero.
Appendix B LMCE Calculation
LMCE is calculated as a linear sensitivity around a DCOPF solution, determining the effect on generation (and therefore emissions) of a single-unit change in load. For ElectricityEmissions.jl, we use the general method set forth in (Lindberg et al., 2021a) to calculate this sensitivity. However, we make several small modifications in order to deal with new constraints introduced by the piecewise-linear cost functions used in our DCOPF formulation.
The DCOPF formulation, as described in Appendix A, has $n=N+2N_{G}$ decision variables, where $N$ is the number of buses and $N_{G}$ is the number of generators. For each bus, the voltage angle $\Theta_{i}$ must be determined, and for each generator $j$, the active power generation $p_{G,j}$ and cost $C_{G,j}$ must be determined.
In the non-degenerate case, this means that the following linear system can be formed

### v3

@@ 4.1. Carbon Intensity Calculation
The first part of the code example demonstrates how to compute the carbon intensity metrics. The MATPOWER-format test-case, containing network data and generator emissions information, is loaded using the parse_file function from the PowerModels.jl package (Coffrin et al., 2018). It is then passed to the functions calculate_ACE, calculate_LMCE, calculate_LACE and calculate_ALMCE to calculate the respective carbon intensity metrics. Each of these functions solve a DC-OPF problem to dispatch the system. Then load demand, optimal generator set-points, and generator carbon intensity (and for some metrics, the network structure) are used to calculate the nodal carbon intensity for each node in the network, as described in Section 3. We calculate total system emissions with the function calculate_system_emissions, which only needs generator set-points as an input.
@@ 4.3. Example Workflow
0. Test-Case Setup First, obtain a MATPOWER-format grid test-case. Test-cases with various properties (size, connectedness, distribution vs. transmission system, etc.), are included as part of the MATPOWER software package (Zimmerman et al., 2011) or are easily accessible via a library maintained by the IEEE Power and Energy Society (Carleton Coffrin, 2024). Second, as shown in Code Block 1, use the PowerModels.jl (Coffrin et al., 2018) package to load the test-case. Third, if not already included in the test-case, assign generator emissions intensities. ElectricityEmissions.jl expects this information for each generator $i$ under the key ”emissions”, e.g. network["gen"][i]["emissions"]. Finally, it is best practice to ensure that no two generators at a given node have the same cost function, as such duplications can lead to problems calculating LMCE/ALMCE (see Appendix A.2 for details). This issue can be easily fixed by adding a small amount of noise to the offending generator cost functions.
@@
2. Solve a Load-Shifting Problem A load uses the carbon intensity values generated in the previous step (or outside carbon metric values/forecasts) as input to solve a load shifting optimization problem to minimize its emissions. Do this for each intensity metric and record the resulting post-shift loads. To estimate the impact of this shifting on the accounted carbon emissions, multiplying the initial carbon intensities by the new load values. Note that these are just estimates, for two main reasons. First, the carbon intensities might change as a result of the load shifting, leading to differences between the estimated and realized accounted carbon emissions. Second, the carbon intensities are themselves estimates of the carbon footprint of electricity. Although a load shifting action leads to a significant reduction in the accounted carbon emissions, it is not guaranteed that the total system emissions would actually decrease by the same amount.
3. Post-Shift Emissions Calculation Shifting load will result in new generator set-points and a new power flow which may change the carbon intensity values at each bus. To determine the final post-shift emissions, we rerun the relevant calculate_...() functions with the shifted load profile to recalculate the carbon intensity metrics. As explained above, the carbon intensities may change as a result of the load shifting. Using the post-shift carbon intensities, it is possible to compute the accounted emissions that would be realized if the load-shifting was implemented. We refer to those as the realized accounted emissions.
@@ 5.1. Dataset
We make use of the RTS-GMLC (Barrows et al., 2020) test case, a popular synthetic dataset geographically “located” in the Southwestern United States. The system contains 73 buses, 120 lines, and 158 generators of varying types, including renewables (wind, solar, and hydro), natural gas, oil, and coal. One year’s worth of hourly load and generation data is available for this system, totaling 8,784 cases. We use the day-ahead load and renewable energy data from the RTS-GMLC data to calculate these metrics on an hourly basis. We make two modifications/additions, similar to the changes made to the same system in (Lindberg et al., 2021a). First, we add data-centers to the system by setting buses 103, 107, 204, and 322 to be data-centers, and refer to those data centers as DC 103, DC 107, DC 204 and DC 322, respectively. Each data center has a nominal load of 250MW, such that the total data-center load corresponds to slightly over 20% of total energy consumption over the course of the year. Second, we use values from the U.S. Department of Energy(Office of Energy Policy and Systems Analysis, U.S. Department of Energy, 2016) to assign the following emissions intensities (in metric tons of CO2 per MWh) to each generator in the system:
@@
Note that the emission values in this table differ from the ones listed in (Lindberg et al., 2021a) (where the emissions values for coal and natural gas appear to have been switched). The data and code are available as examples in the package documentation.  [^3] https://github.com/WISPO-POP/ElectricityEmissions.jl
@@ 5.3. Results
Generated System DC Total DC 103 DC 107 DC 204 DC 322
Sum Sum Sum Sum Sum Sum Sum
LMCE 15.828 33.012 6.692 1.686 1.562 1.917 1.527
ALMCE 15.828 3.162 0.803 0.679 1.035 0.644
ACE 15.828 3.008 0.752 0.752 0.752 0.752
LACE 15.828 2.707 0.577 0.850 1.153 0.126
Table 1. Whole-Year Emissions Accounting (Million Tons CO2/MWh)
System DC 103 DC 107 DC 204 DC 322
Mean SD Mean SD Mean SD Mean SD Mean SD
@@
Table 2. Whole-Year Emissions Intensity (Tons CO2/MWh)
@@
We first consider how the choice of carbon intensity metric impacts the total accounted carbon emissions across the year. Table 1 contains the total system emissions as measured by the generator set-points as well as the accounted emissions for the entire system, the total data center load (denoted by DC load) and for each individual data-center (denoted by DC followed by the corresponding load bus index). In looking at the emissions assigned to all loads in the system, we see that ALMCE, ACE, and LACE lead to accounted emissions that equal the generated emissions value of 15.8 million tons of CO2. LMCE, however, lead to much higher accounted emissions at 33.0 million tons. The trend of much higher emissions with LMCE also holds true for the different data center locations.
@@
We next examine the assigned emissions intensities for each carbon metric. Table 2 shows summary statistics of emissions intensity across the all nodes in the system, and for each data-center load individually. We show both the mean value across all hours of the year, as well as the standard deviation of the emissions intensity. First, we analyze inter-metric intensity differences for the system as a whole. On average, LMCE has the highest intensity values across the whole system, with a mean value of 0.740 Tons CO2/MWh. This is dramatically higher than the next-highest, ACE, with a mean intensity of 0.342 Tons CO2/MWh. ACE shows the least whole-system variability out of all the metrics, likely due to the fact that it does not vary spatially within a given time-step.
Second, we analyze emissions intensities across the four data-centers. As in the whole-system statistics, LMCE assigns the highest average intensity values of any metric to all data-centers. However, the relative magnitude of ALMCE, ACE, and LACE varies greatly between location. Out of these three metrics, LACE has the highest mean value for the DC 107 and DC 204, but the lowest for DC 103 and DC 322. Despite this variability, all of the metrics (save ACE, which is location-agnostic) are in agreement on the ranking of average emissions intensity between data-center locations. DC 204 has the highest emissions intensity, followed by DC 103, DC 107, and finally DC 322.
@@ 6.1. Dataset
We make use of the same modified RTS-GMLC test case as described in the accounting-only case study, which provides a year’s worth of hourly load and generation data. Whereas the accounting-only case study added fixed 250MW data-center loads, the data-centers are new allowed to increase or decrease their load by up to 20% relative to the 250MW baseline in order to minimize emissions.
@@ 6.2. Methodology
We perform load shifting based on LMCE, ALMCE, ACE, and LACE carbon metrics, within each 24-hour period of year, following the workflow described in Section 4.3 (note that the test-case setup is described in the paragraph above.
1. Initial Emissions Calculations We first calculate initial ”pre-shift” emissions intensity for each metric, as would be determined from a DA market clearing solution. This provides a forecast of the carbon metric for each node, and at each hour of the day.
2. Solve Data Center Load Shifting Problem Next, we optimize the shifting of electricity consumption between the data-centers, which are assumed to be controlled by a single entity. We assume that the data center operator can choose where and when certain computing workloads are being processed. This ability to shift workloads translates into an ability to adjust electricity consumption across locations and time steps. Specifically, the data center operator can shift workloads and associated electricity consumption to align with periods and locations with low carbon intensity, thus minimizing carbon emissions. We assume that the data centers are able to increase or decrease their electricity consumption by up to 20% relative to their average consumption (which is set to 250MW in our case) at any given time step, but that the total energy consumption across all data centers for one day (a surrogate for the total computational work that is done) has to remain the same. The specific mathematical formulation of the data center load shifting problem can be found in Appendix A.3, along with some notes on the potential impact of parameter choice and case-specific modeling decisions. We note that in our case study, we assume that only the data centers shift their load, though the simulation setup could be extended to handle (independent) load shifting by other loads as well.
3. Post-Shift Emissions Calculation From the solution to the load-shifting optimization problem, we obtain estimated accounted emissions for the data-centers. These estimated accounted emissions are calculated based on the new data-center loading values and the old (pre-shift) carbon intensities. However, the adjustment to the load consumption then affects the actual load profile in the RT energy market. We subsequently solve the OPF for the RT market, and calculate the realized emissions metrics, after the load shifting has occurred.
@@ 6.3. Results
Pre-Shift Accounted
Emissions
Post-Shift Accounted
Emissions (Estimated)
Post-Shift Accounted
Emissions (Realized)
@@
System DC Loads Non-DC Loads DC Loads System DC Loads Non-DC Loads System
@@
Table 3. Shifting Case Study: Emissions Results (Million Tons CO2)
@@
Starting on the left, we show the pre-shift accounted emissions which are the carbon accounting values prior to any shifting (i.e. they are the same values as in the accounting case study, summarized here for convenience of the reader).
Next we show the post-shift accounted emissions, which are divided into ”estimated” and ”realized”. Estimated post-shift accounted emissions, which refer to the data-centers’ belief about what their accounted emissions will be after shifting, based on their new (post-shift) load values and the old (pre-shift) carbon metric values that they were optimizing for. We only show estimated post-shift accounted emissions for the data centers themselves, as these are the only loads that perform load shifting and obtain such estimates. The realized post-shift emissions are the actual accounted emissions values, determined via post-shift loading and post-shift (i.e. re-calculated) carbon intensity values.
@@
We begin by investigating whether load-shifting leads to a reduction in the accounted emissions assigned to data-center loads. To analyze this, we first analyze the realized accounted emissions for DC loads. We observe that shifting according to any of the metrics leads to a decrease in the accounted emissions for DC loads, but to varying degrees. The greatest reduction in accounted emissions was achieved by ALMCE, followed by LACE and ACE. The smallest reduction in accounted emissions happened when the data centers shifted according to LMCE.
@@
In conclusion, these results indicate that even though data centers are able to claim emission reductions after shifting with ALMCE, ACE and LACE, these emission reduction come from reallocating emissions to other loads rather than from actually decreasing total generator emissions. In contrast, when shifting with LMCE, the data centers actually reduced total system emissions.
@@
While over the course of the whole year, LMCE/LACE-based shifting seemed to decrease emissions and ALMCE/ACE-based shifting seemed to increase them, there is a large amount of variability in the performance of each metric on any particular day. Figure 4 shows the distribution of the daily changes in system emissions resulting from shifting with the different metrics, as well as the mean, median, 10th percentile and 90th percentile of the reductions. We can clearly see that shifting based on LMCE is much more likely to lead to large carbon emission reductions (i.e. large negative values) and less likely to increase the total carbon emissions from the overall system (i.e. has fewer positive values). We can also clearly see how some of the other metrics have a tendency to increase system carbon emissions rather than reducing them.
@@
Finally, we briefly examine the outcome of shifting based on one metric and accounting using another (e.g. determining data-center load shifting with LMCE, then assigning emissions using ACE). Figure 5 shows the absolute change in assigned data-center emissions, defined as the difference between pre- and post-shift accounted emissions (measured by the accounting metric).
As is to be expected, when the same metric is used for shifting and accounting, we see the greatest reduction in accounted emissions. There is great variation in the ’cross-metric’ behavior of different ”shifting metric”-”accounting metric” pairs. For many of them, we see that shifting with one metric (e.g. LMCE) increases the accounted carbon emissions when a different metric is used for accounting. This indicates that companies or organizations who want to not only reduce overall carbon emissions, but also the carbon emissions that are allocated to them, have strong incentives to shift according to the same metric that is used for accounting.
@@ 7. Conclusions
In our case study, we find that load shifting based on the ACE, ALMCE and LACE intensity metrics has very little impact on overall carbon emission and instead tends to shift emissions from the data centers (which actively seek to minimize their emissions according to these metrics) to non-data center loads. In contrast, shifting according to LMCE reduces emissions for both data centers, non-data center loads, and the overall system. However, it is challenging to do accounting with LMCE, as the allocated carbon emissions are more than two times higher than the actual carbon emissions of the system.
@@
In future work, we plan to expand the package to include other metrics and other functionality, allowing input from the community. In particular, we want to add support quadratic generator cost functions and alternative OPF formulations. We also wish to design new metrics with nice properties for both accounting and load shifting, and perform more thorough analysis of the differences between metrics on more realistic power system test cases and more realistic load-shifting setups.
Appendix A Appendix
A.1. DCOPF Formulation
We use the following DCOPF formulation with piecewise linear generator cost to dispatch generation in ElectricityEmissions.jl, and for the linear generation sensitivity calculations necessary for the LMCE and ALMCE carbon metrics:
@@ A.1. DCOPF Formulation
Each generator $j$ has a piecewise-linear cost function with $N_{ct}$ terms. Equation (6b) constrains the cost of each generator $j$ to be above all of the $k$ linear functions that define its segments, with $a_{j,k}$ and $b_{jk}$ representing the slope and intercept of the function. Equation (6c) is the standard nodal power conservation constraint, referencing the subset of generators $G_{i}\subset\mathcal{G}$, loads $D_{i}\subset\mathcal{D}$, and lines $(i,j)\in L$ connected at bus $i$. Equation (6d) limits the flow on each line $(i,j)$ to be below the maximum capacity $P_{i,j}^{lim}$, and Equation (6e) sets limits for each generator $j$ to be between $P_{G,j}^{min}$ and $P_{G,j}^{max}$. Finally, Equation (6f) sets the voltage angle of the reference bus to zero.
A.2. LMCE Calculation
LMCE is calculated using linear sensitivity analysis for the DCOPF solution, and describes the effect on generation and emissions due to a small change in load. For ElectricityEmissions.jl, we use the general method set forth in (Lindberg et al., 2021a) to calculate this sensitivity. However, we make several small modifications in order to deal with the new constraints and variables introduced by the piecewise-linear cost functions used in our DCOPF formulation (where (Lindberg et al., 2021a) simply assumed linear costs).
The DCOPF formulation, as described in Appendix A.1, has $n=N+2N_{G}$ decision variables, where $N$ is the number of buses and $N_{G}$ is the number of generators. For each bus, the voltage angle $\Theta_{i}$ must be determined, and for each generator $j$, the active power generation $p_{G,j}$ and cost $C_{G,j}$ must be determined.
For any instance of the DCOPF problem, which is a linear program, there exists at least on basic optimal solution. At this solution, we can identify $n$ linearly independent constraints that are being satisfied with equality. For this solution, the following linear system can be formed